Sharu has $2.35 in nickels and dimes. If he has a total of thirty-two coins, how many of each coin does he have?

Answer by drk(1908)

This is a classic money problem. There are 2 equations - first the value equation and second the number equation. Here they are:
(i) 5n + 10d = 235
(ii) n + d = 32
Multiply the second equation by -5.
(i) 5n + 10d = 235
(iii)-5n -5d = -160
Add (i) and (iii) together. We get
5d = 75.
Solve for d.
d = 15.
Now find n.
n = 17.